a beaker contains 15 milliliters of 20% alcohol solution. how many milliliters of a 60% alcohol solution should be added in order to make a 30% solution?
Answer by checkley77(12844)

.20*15+.60x=.30(15+x)
3+.60x=4.5+.30x
.60x-.30x=4.5-3
.30x=1.5
x=1.5/.30
x=5 ml. of 60% alcohol is used.
Proof:
.20*15+.60*5=.3(15+5)
3+3=.3*20
6=6